OpenStudy (anonymous):

What is the sum of the infinite geometric series 1/2 +1/4 + 1/8 + 1/16 +... ?

OpenStudy (anonymous):

well, you can see the pattern, right?

OpenStudy (anonymous):

a1=1/a0, a2=1/2*a1, a3=1/2*a2, etc.

OpenStudy (anonymous):



OpenStudy (anonymous):

converges to 1

OpenStudy (anonymous):

there is an exact formula but I don't remember sorry...

OpenStudy (anonymous):

\[\sum_{n=0}^{\infty}\frac{ 1 }{ 2n }\]

OpenStudy (anonymous):

I think that's about right.

OpenStudy (anonymous):

it's geometric.

OpenStudy (anonymous):

the formula is Sn=(t1(1-r^n))/1-r

jimthompson5910 (jim_thompson5910):

Any infinite geometric series has an infinite sum of S = a/(1-r) where a = first term r = common ratio and |r| < 1 If |r| > 1, then the series doesn't converge

OpenStudy (anonymous):

the formula for the sum uses a=first term r=ration you're multiplying by and the formula is a/(1-r) if I remember correctly

OpenStudy (anonymous):

whats the n amount? in the formula i listed above?

jimthompson5910 (jim_thompson5910):

the formula you listed sparky16 is only used for partial sums

OpenStudy (anonymous):

oh is it this formula: S=t1/1-r and r has to be less than 1

jimthompson5910 (jim_thompson5910):

more like |r|<1, but yes

OpenStudy (anonymous):

so after plugging that in, I got 1. Is that correct?

jimthompson5910 (jim_thompson5910):

yes S = a/(1-r) S = (1/2)/(1-1/2) S = (1/2)/(1/2) S = 1

OpenStudy (anonymous):

thank you so much(:

jimthompson5910 (jim_thompson5910):

np